### v2

# Byung-Kwan Lee KAIST

Abstract
Large language and vision models (LLVMs) have been driven by the generalization power of large language models (LLMs) and the advent of visual instruction tuning. Along with scaling them up directly, these models enable LLVMs to showcase powerful vision language (VL) performances by covering diverse tasks via natural language instructions. However, existing open-source LLVMs that perform comparably to closed-source LLVMs such as GPT-4V are often considered too large (e.g., 26B, 34B, and 110B parameters), having a larger number of layers. These large models demand costly, high-end resources for both training and inference. To address this issue, we present a new efficient LLVM family with 1.8B, 3.8B, and 7B LLM model sizes, Traversal of Layers ([Uncaptioned image]TroL), which enables the reuse of layers in a token-wise manner. This layer traversing technique simulates the effect of looking back and retracing the answering stream while increasing the number of forward propagation layers without physically adding more layers. We demonstrate that [Uncaptioned image]TroL employs a simple layer traversing approach yet efficiently outperforms the open-source LLVMs with larger model sizes and rivals the performances of the closed-source LLVMs with substantial sizes. Code is available in https://github.com/ByungKwanLee/TroL.
TroL: Traversal of Layers for Large Language and Vision Models
Byung-Kwan Lee KAIST leebk@kaist.ac.kr                        Sangyun Chung KAIST jelarum@kaist.ac.kr                        Chae Won Kim KAIST chaewonkim@kaist.ac.kr
Beomchan Park KAIST bpark0810@kaist.ac.kr                        Yong Man Ro KAIST ymro@kaist.ac.kr
1 Introduction
Figure 1: Overview of Refer to captionTroL’s layer traversing
Figure 2: Performances for efficient LLVM family, Refer to captionTroL, across three model sizes: 1.8B, 3.8B, and 7B
The great success of closed-source large language and vision models (LLVMs) such as GPT-4V (Achiam et al., 2023), Gemini-Pro (Team et al., 2023), and Qwen-VL-Plus (Bai et al., 2023) has prompted the world to build open-source LLVMs using open-source large language models (LLMs) and visual instruction tuning (Liu et al., 2023c, b, 2024a). These movements are expected to contribute significantly to both research and industry through numerous downstream tasks, such as on-device chatbot systems.
In order to bootstrap their vision language performance, several studies have curated high-quality visual instruction tuning datasets (Chen et al., 2023b, 2024b) by leveraging the power of closed-source LLVMs. Further, they have physically increased the model sizes (McKinzie et al., 2024; Li et al., 2024b; Liu et al., 2024a) to enhance their capability to understand complex question-answer pairs.
Following these efforts, there has been a recent rise in techniques for employing additional modules: mixed vision encoders (Kar et al., 2024; Lu et al., 2024; Goncharova et al., 2024; Ranzinger et al., 2023; Zhao et al., 2024), multiple computer vision models (Chen et al., 2024a; Wang et al., 2024b; Jiao et al., 2024; Lee et al., 2024b, c), and mixtures of experts (MoE) for efficiently scaling LLVMs to fulfill specific purposes (Lin et al., 2024; Lee et al., 2024c; Li et al., 2024a, c; Guo et al., 2024; McKinzie et al., 2024; Li et al., 2024b; Gao et al., 2024; Sun et al., 2024a). In the end, LLVMs combined with these techniques have shown improved vision language performances, achieving results comparable to directly scaled-up LLVMs with 26B, 34B, and 110B model sizes. Additionally, some approaches have demonstrated performances that surpass those of the closed-source LLVMs.
Figure 3: Overview of vision language performances compared with the Refer to captionTroL and closed-source LLVMs
However, directly scaling the model size up or using additional modules may not be considered a fundamental solution to enlarging learning capabilities regarding complex question-answer pairs. This is because they physically add a considerable number of training parameters or borrow richer knowledge from external modules. In other words, it remains unexplored how LLVMs with smaller model sizes can effectively enhance learning capabilities despite their inherent physical limitations.
To address this, simply increasing the image resolution size (Li et al., 2023a; Bai et al., 2023; Wang et al., 2023; Ye et al., 2023b; Hu et al., 2024a) and dynamically dividing images into sub-parts for hierarchical focus (Liu et al., 2024a; McKinzie et al., 2024; Xu et al., 2024) may be good candidates to achieve our purpose without employing any additional modules mentioned. Nonetheless, these strategies are mostly intended to embed rich image information to further improve overall and fine-grained image understanding. Hence, we need to focus on how multimodal LLMs can be enhanced by themselves. Once more efficient LLMs that perform comparably to closed-source models are released publicly, it will mitigate the necessity for high-end GPUs and accelerate significant advancements in various downstream applications, including on-device processing.
Therefore, we present a new efficient LLVM family with 1.8B, 3.8B, and 7B model sizes, Traversal of Layers ([Uncaptioned image]TroL), which enables the reuse of layers in a token-wise manner. To overcome the inherent limitations of smaller-sized LLVMs, we opt to increase the number of forward propagations rather than physically adding more layers, as is normally done in scaled-up LLVMs. This technique, which we call layer traversing, allows LLVMs to retrace and re-examine the answering stream, akin to human retrospection and careful thought before responding with an answer. Figure 1 represents how the layer traversing technique is practically implemented in TroL-Layer, where TroL-Mixer serves as the token-wise mixing operation under lightweight additional parameters: 49K, 98K, and 131K in total layers. This is a significantly tiny number compared with the 1.8B, 3.8B, and 7B model sizes.
To successfully apply layer traversing to LLVMs, we employ a two-step training process, establishing [Uncaptioned image]TroL. The first step involves training a vision projector and all TroL-Mixers for each TroL-Layer. This is a crucial step because it not only aligns vision and language information but also tunes the TroL-Mixers with the answering stream in backbone multimodal LLMs, thereby facilitating the use of layer traversing. The second training step includes further training of these components along with the backbone multimodal LLMs. To achieve efficient training, we use Q-LoRA (Dettmers et al., 2023) training for the backbone multimodal LLMs under 4/8-bit quantization.
In conducting the two-step training, we demonstrate that [Uncaptioned image]TroL is an efficient model, yet it outperforms open-source LLVMs with larger model sizes (e.g., 26B, 34B, 72B, and 110B) and closed-source LLVMs with a substantially vast amount of parameters, as illustrated in Figures 2 and 3.
Our contribution can be summarized into two main aspects:
  • We introduce a new efficient LLVM family—1.8B, 3.8B, and 7B, Traversal of Layers ([Uncaptioned image]TroL), which enables the reuse of layers, simulating the effect of retracing the answering stream.
  • [Uncaptioned image]
    TroL proves its superior effectiveness on various evaluation benchmarks compared with substantially sized open- and closed-source LLVMs without directly scaling up the model size and without any additional modules.
2 Related Works
Figure 4: Overview of two-step training to build an efficient LLVM Family, Refer to captionTroL
Large Language and Vision Models.
The curation of visual instruction tuning datasets  (Liu et al., 2023c, b, 2024a; Dai et al., 2023) has significantly propelled the rapid development of LLVMs (Chen et al., 2023a; Bai et al., 2023; Zhu et al., 2023; Li et al., 2023b; Ye et al., 2023a, b; Chen et al., 2023b; Contributors, 2023; Zhang et al., 2023; Chen et al., 2023c, 2024d). Building upon this foundation, recent research has further advanced LLVMs’ capabilities through two key strategies: scaling up model sizes and designing specialized instruction tuning datasets. Firstly, increasing model sizes has emerged as a prominent approach to enhancing LLVM performance and capacity. For example, LLVMs (McKinzie et al., 2024; Li et al., 2024b; Liu et al., 2024a; Wang et al., 2023; Laurençon et al., 2023; Sun et al., 2023; Gao et al., 2024; Sun et al., 2024a) deploy larger architectures and more parameters to enlarge the representational power of these models. Moreover, the development of meticulous instruction tuning datasets (Chen et al., 2023b; Li et al., 2024b; Hu et al., 2024a; Gao et al., 2023; Wang et al., 2024a; Yue et al., 2023, 2024) has played a pivotal role in improving LLVMs for specific tasks or domains.
Additionally, recent research has explored more direct approaches to enhance LLVM’s image perception capabilities, either by modifying visual input and utilizing additional modules. For example, Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023), and mPLUG family (Ye et al., 2023b; Hu et al., 2024a) increase image resolution to enrich visual information in LLVMs. Moreover, few approaches (Liu et al., 2024a; McKinzie et al., 2024; Li et al., 2024b) improve visual tokens through image partitioning. Additionally, the integration of additional vision encoders (Kar et al., 2024; Lu et al., 2024; Goncharova et al., 2024; Ranzinger et al., 2023; Zhao et al., 2024) and external computer vision modules (Chen et al., 2024a; Wang et al., 2024b; Jiao et al., 2024; Lee et al., 2024b, c) augment LLVMs’ image perception capability, thereby enhancing overall performance on multimodal tasks.
While these lines of works expand the overall learning capabilities of LLVMs, they do not necessarily enhance the fundamental capacity of LLVMs. Consequently, there remains a need for further research into the intrinsic mechanisms of LLVMs without scaling models or leveraging additional modules. Thus, we introduce a new LLVM family, [Uncaptioned image]TroL, specifically tailored to enhance the learning capabilities of LLVMs, where layer traversing is presented to efficiently reuse layers. This approach simulates the effect of retracing the answering stream, offering a focused solution to propel the advancement of LLVMs.
Numerous Efficient Approaches.
Despite the remarkable achievements of LLVMs in a short period of time, closed-source LLVMs require a significant number of parameters and resources. For that reason, many LLVMs have been exploring methods to create more efficient models. These methods aim to reduce the number of parameters or improve computational efficiency without significantly compromising performance. Numerous methods to reducing the number of parameters involve sharing specific weights within the model (Thawakar et al., 2024; Lan et al., 2020; Takase and Kiyono, 2023; Reid et al., 2021), eliminating weights that contribute less to the performance (Sun et al., 2024b; Ma et al., 2023; Cao et al., 2023; Frantar and Alistarh, 2023; Men et al., 2024), and employing quantization techniques (Shao et al., 2024a; Li et al., 2023f; Park et al., 2024a).
These efforts primarily focus on reducing inference time or accelerating training speed while maintaining performance, rather than fundamentally improving it. In contrast, [Uncaptioned image]TroL aims to efficiently enhance the learning capabilities of LLVMs in understanding complex question-answer pairs.
[…]
BLIP2-13B Li et al. (2023d) - 61.0 - - 46.4 85.3 - 1584 - - - 22.4 -
InstructBLIP-7B Dai et al. (2023) 56.7 60.5 - - 53.4 - 53.6 - 25.3 36.0 23.9 26.2 -
InstructBLIP-13B Dai et al. (2023) - 63.1 - - - 78.9 - - - 33.9 - 25.6 -
IDEFICS-9B Laurençon et al. (2023) 51.5 - - - - 74.6 - 1353 19.8 48.2 25.2 23.7 -
Qwen-VL-7B Bai et al. (2023) 59.4 67.1 - - - - - - - 38.2 7.4 - -
Qwen-VL-Chat-7B Bai et al. (2023) 33.8 68.2 - - 58.2 - 56.4 1849 - 60.6 56.7 47.3 -
MiniGPT-4-7B Zhu et al. (2023) 51.8 - - - - - - - 23.1 23.0 11.9 22.1 -
Otter-7B Li et al. (2023b) 47.2 - - - - 72.5 - 1599 19.7 48.3 - 24.7 -
UIO-2-XXL-6.8B Lu et al. (2023a) - 86.2 - - 61.8 87.7 - - - 71.5 - - -
LLaVA-7B Liu et al. (2023c) - 38.5 - - - 80.2 44.1 1055 - 34.1 14.1 26.7 -
LLaVA1.5-7B Liu et al. (2023b) 60.1 66.8 - - 58.6 85.9 - 1805 - 64.3 58.3 30.5 63.4
LLaVA1.5-13B Liu et al. (2023b) 61.4 71.6 54.8 18.2 61.6 85.9 46.7 1826 27.6 67.7 63.6 35.4 -
mPLUG-Owl-7B Ye et al. (2023a) 58.9 - - - - - - - 22.2 46.6 - - -
mPLUG-Owl2-7B Ye et al. (2023b) 62.9 68.7 - - - - - - 64.5 60.3 36.2 -
ShareGPT4V-7B Chen et al. (2023b) 63.4 68.4 - - 69.7 - 49.8 1944 25.8 68.8 62.2 37.6 -
InternLM-XC-7B Zhang et al. (2023) 64.4 - - - 66.1 - 57.0 1919 29.5 74.4 72.4 35.2 -
Monkey-10B Li et al. (2023g) - 69.4 - - 68.9 - 58.4 1924 34.8 72.4 67.5 33.0 -
VILA-7B Lin et al. (2023b) - 68.2 - - 61.1 85.5 - - - 68.9 61.7 34.9 -
VILA-13B Lin et al. (2023b) - 73.7 - - 62.8 84.2 - - - 70.3 64.3 38.8 -
SPHINX-7B Lin et al. (2023c) - 70.6 - - 71.6 86.9 - 1797 27.8 65.9 57.9 40.2 -
SPHINX-MoE-7B×$\times$×Gao et al. (2024) 66.2 70.6 - - 73.0 89.6 - 1852 42.7 71.3 - 40.9 -
SPHINX-Plus-13B Gao et al. (2024) 66.2 70.6 - - 74.8 89.1 52.1 1741 36.8 71.0 - 47.9 -
LLaVA-NeXT-7B Liu et al. (2024a) - 70.1 - - 70.2 86.5 - 1851 34.6 69.6 63.3 43.9 72.3
LLaVA-NeXT-8B Liu et al. (2024a) - - 71.6 69.5 - - - 1972 37.5 72.1 - - 80.1
LLaVA-NeXT-13B Liu et al. (2024a) - 73.6 70.0 62.2 72.2 86.7 - 1892 35.1 70.0 68.5 47.3 72.3
MM1-7B McKinzie et al. (2024) - 72.6 - - 69.9 86.6 - 1858 35.9 72.3 - 42.1 -
MM1-MoE-7B×$\times$32 McKinzie et al. (2024) - 74.4 - - 70.9 87.8 - 1992 40.9 72.7 - 45.2 -
MiniGemini-HD-7B Li et al. (2024b) - - - - - - - 1865 32.2 65.8 - 41.3 -
MiniGemini-HD-13B Li et al. (2024b) - - - - - - - 1917 37.0 68.6 - 50.5 -
TroL-7B 73.6 92.8 78.5 71.2 75.3 87.8 65.3 2308 51.8 83.5 81.2 54.7 92.8
[…]
3 [Uncaptioned image]TroL: Traversal of Layers
Model Architecture.
As illustrated in Figure 4, [Uncaptioned image]TroL is composed of a vision encoder, a vision projector, and a backbone multimodal large language model (MLLM) based on a pre-trained LLM. We utilize CLIP-L (Radford et al., 2021) and InternViT (Chen et al., 2023c) for the vision encoder, which are text-aligned vision encoders based on image-text contrastive learning with a small text encoder (CLIP) and QLLaMA-8B (Cui et al., 2023) (InternViT), respectively. For the vision projector, we employ two fully-connected layers with the GELU activation function (Hendrycks and Gimpel, 2016). As for the backbone multimodal LLM, we use Phi-3-mini (Abdin et al., 2024) with a 3.8B model size, and InternLM2 (Team, 2023; Cai et al., 2024) with 1.8B and 7B model sizes. 3.3T and 2T tokens are used during the pre-training of these LLMs, respectively.
Visual Instruction Tuning Dataset.
We gather a wide range of visual instruction tuning datasets requiring diverse capabilities such as fundamental image understanding, common-sense knowledge, non-object concepts (e.g., charts, diagrams, documents, signs, symbols), math problems, and their integrated capabilities. This is because we aim to make [Uncaptioned image]TroL encompass diverse capabilities for vision language tasks despite its efficient model size. To balance the dataset samples across numerous capabilities, we selectively choose samples from existing visual instruction tuning datasets: ShareGPT4V-Caption/Instruct (Chen et al., 2023b), ALLaV4V-Text (Chen et al., 2024b), MiniGemini-Instruct (Li et al., 2024b), Doc-Downstream/Reason (Hu et al., 2024a), GLLaVA-Align/Instruct (Gao et al., 2023), and Math-Vision/Instruct/Plus (Wang et al., 2024a; Yue et al., 2023, 2024). In summary, we collect 899K real-world image/text-only samples, 627K samples for documents, charts, diagrams, signs, and symbols, and 747K math samples (180.5K with images and 566.8K text-only). Overall, the number of visual instruction tuning samples we used to build [Uncaptioned image]TroL totals 2.3M samples.
[…]
UIO-2-XL-3.2B Lu et al. (2023a) - 87.4 - - 60.2 87.2 - - - 68.1 - - -
Gemini Nano-2-3.2B Team et al. (2023) - - - - - - - - 30.6 - - - -
MobileVLM-3B Chu et al. (2023) - 61.2 - - - 84.9 - - - 59.6 - - -
MobileVLM-V2-3B Chu et al. (2024) - 70.0 - - - 84.7 - - - 63.2 - - -
MoE-LLaVA-2.7B×$\times$×Lin et al. (2024) - 70.3 - - - 85.7 - - - 68.0 - 35.9 -
LLaVA-Phi-2.7B Zhu et al. (2024) - 68.4 - - - 85.0 - - - 59.8 - 28.9 -
Imp-v1-3B Shao et al. (2024b) - 70.0 - - - 88.0 - - - 66.5 - 33.1 -
TinyLLaVA-3.1B Zhou et al. (2024) - 69.1 - - - 86.4 - - - 66.9 - 32.0 -
TinyLLaVA-Sig-Phi-3.1B Zhou et al. (2024) - 69.1 - - - 86.4 - - - 66.9 - 32.0 -
Bunny-3B He et al. (2024) - 70.9 38.2 - 62.5 86.8 - 1778 - 68.6 - - -
MiniCPM-2.4B Hu et al. (2024b) - - 56.3 - - - - 1650 28.9 64.1 62.6 31.1 -
MiniCPM-V2-2.8B Hu et al. (2024b) - - 62.9 - - - - 1809 38.7 69.1 66.5 41.0 -
MM1-3B McKinzie et al. (2024) - 69.4 - - 68.8 87.4 - 1762 32.0 67.8 - 43.7 -
MM1-MoE-3B×$\times$64 McKinzie et al. (2024) - 76.1 - - 69.4 87.6 - 1773 32.6 70.8 - 42.2 -
ALLaVA-3B Chen et al. (2024b) - - - - 65.2 - - 1623 - 64.0 - 32.2 -
ALLaVA-3B-Longer Chen et al. (2024b) - - - - 65.6 - - 1564 - 64.6 - 35.5 -
TroL-3.8B 70.0 90.8 73.6 73.8 70.5 86.5 62.2 1980 55.1 79.2 77.1 51.1 76.6
UIO-2-L-1.1B Lu et al. (2023a) - 78.6 - - 51.1 77.8 - - - 62.1 - - -
MobileVLM-1.7B Chu et al. (2023) - 57.3 - - - 84.5 - - - 53.2 - - -
MobileVLM-V2-1.7B Chu et al. (2024) - 66.7 - - - 84.3 - - - 57.7 - - -
DeepSeek-VL-1.3B Lu et al. (2024) - - - - 66.7 87.6 - - 31.1 64.6 62.9 34.8 -
Mini-Gemini-2B Li et al. (2024b) - - - - - - - 1653 29.4 59.8 - - -
MoE-LLaVA-1.8B×$\times$×Lin et al. (2024) - 63.1 - - - 87.0 - - - 59.7 - 25.3 -
TroL-1.8B 68.2 87.5 68.9 64.0 69.0 88.6 60.1 2038 45.4 76.1 74.1 45.1 69.7
[…]
Figure 5: Overview of mix operation in TroL-Mixer
Layer Traversing.
To effectively enlarge learning capabilities with smaller sized LLVMs, we introduce a layer traversing technique that allows the reuse of layers. As described in Figure 1, once the input token $x\in\mathbb{R}^{N\times D}$ (i.e., vision language features), where $N$ denotes the number of tokens and $D$ denotes the hidden dimension of a layer, is given, then a layer normally outputs $L(x)$ from the input token. Layer traversing makes the output of the layer forward in the equal layer once again: $L(L(x))$. Subsequently, the outputs $L(x)$ and $L(L(x))$ from the equal layer get mixed to further improve the vision language features by themselves. Here, we present TroL-Mixer that provides a mixing operation with $L(x)$ and $L(L(x))$, where TroL Gating is introduced to determine how much reused vision language feature $L(L(x))$ is needed for the next layer, by looking at the feature status of the first propagation output $L(x)$. More specifically, the output $L(x)$ is propagated into the TroL Gating and it produces a mixing ratio $w\in\mathbb{R}^{N}$ for each token. It is used for token-wise multiplication $\odot$ with $L(x)$ and $L(L(x))$. Finally, the mixed output $(1-w)\odot L(x)+w\odot L(L(x))$ in Figure 5 is used to propagate to the next layer and it is continually operated as it goes layers. By doing this, we expect the layer traversing to stimulate the effect of retracing and looking back at the answering stream once again, thereby enlarging the learning capabilities by nature.
Training Strategy.
TroL-Layer is applied to backbone multimodal LLMs described in Figure 1 and 4 for application to layer traversing technique. Thereafter, we conduct a two-step training process to effectively implement layer traversing using LLVMs, creating a new efficient LLVM family named [Uncaptioned image]TroL. In the first training step, we focus on training a vision projector and all TroL-Mixers for every TroL-Layer. This step is essential as it aligns vision and language information while synchronizing the TroL-Mixers with the response stream in the backbone multimodal LLMs, thus facilitating the understanding of layer traversing operation. The subsequent second training step involves additional training of these elements alongside the backbone multimodal LLMs together.
4 Experiment
Benchmarks OmniFusion-7B DeepSeek-VL-7B MoVA-7B ASMv2-7B LAF-7B CoLLaVO-7B MoAI-7B TroL-1.8B TroL-3.8B TroL-7B
POPE 87.2 88.1 88.6 86.3 88.8 87.2 87.1 88.6 86.5 87.8
[…]
(a) Comparison with LLVMs using additional modules: OmniFusion Goncharova et al. (2024), DeepSeek-VL Lu et al. (2024), MoVA Kar et al. (2024), ASMv2 Wang et al. (2024b), LAF Jiao et al. (2024), CoLLaVO Lee et al. (2024b), and MoAI Lee et al. (2024c)
LLVMs Recognition OCR Knowledge Language Generation Spatial Awareness Math Problems Avg
CoLLaVO-7B Lee et al. (2024b) 45.6 31.1 29.8 30.2 37.9 5.8 41.0
[…]
TroL-1.8B 42.0 48.2 31.9 29.0 47.1 41.2 45.1
TroL-3.8B 45.7 56.6 37.0 40.6 56.5 48.1 51.1
TroL-7B 54.2 54.6 42.4 49.3 52.7 53.8 54.7
(b) Evaluating sub-benchmark in MM-Vet Yu et al. (2023) with LLVMs utilizing computer vision models
LLVMs CP FP IR LR ST MA Avg
Yi-VL-34B Young et al. (2024) 53.2 31.2 52.0 32.4 12.4 35.2 36.1
[…]
TroL-1.8B 63.2 41.6 59.2 47.2 30.0 31.6 45.5
TroL-3.8B 61.2 40.0 54.0 43.2 31.2 49.6 46.5
TroL-7B 69.2 47.6 65.6 45.2 37.2 42.8 51.3
(c) MMStar Chen et al. (2024c)
LLVMs TD TL TO VI VD VO Avg
G-LLaVA-7B Gao et al. (2023) 20.9 20.7 21.1 17.2 16.4 9.4 16.6
[…]
TroL-1.8B 26.1 26.5 25.5 25.6 25.6 14.8 24.0
TroL-3.8B 42.3 38.8 40.6 35.5 35.9 21.4 35.8
TroL-7B 37.8 34.1 36.9 32.1 32.1 19.5 32.1
(d) MathVerse Zhang et al. (2024)
LLVMs Website Element Action Average
Caption WebQA HeadOCR OCR Ground Prediction Ground
InstrcutBLIP-13B Dai et al. (2023) 11.6 5.2 7.6 6.0 11.4 11.4 17.5 10.1
Yi-VL-6B Young et al. (2024) 8.0 14.3 43.8 3.5 16.2 13.9 13.6 16.2
LLaVA1.5-7B Liu et al. (2023b) 15.3 13.2 41.0 5.7 12.1 17.8 13.6 17.0
LLaVA1.5-13B  Liu et al. (2023b) 20.0 16.2 41.1 11.8 15.0 22.8 8.7 19.4
CogVLM-17B Wang et al. (2023) 16.6 30.6 65.9 10.0 17.7 11.7 23.3 25.1
VILA-13B Lin et al. (2023b) 12.7 28.8 67.9 12.6 16.5 36.3 16.5 27.3
DeepSeek-VL-7B Lu et al. (2024) 18.1 30.0 63.4 18.1 16.2 35.2 15.5 28.1
LLaVA-NeXT-7B Liu et al. (2024a) 27.0 39.8 57.3 54.8 31.7 30.6 10.7 36.0
LLaVA-NeXT-13B Liu et al. (2024a) 26.5 44.5 52.8 56.1 31.7 48.4 15.5 39.4
LLaVA-NeXT-34B Liu et al. (2024a) 24.3 48.2 67.1 71.9 43.1 74.0 25.2 50.5
TroL-1.8B 14.2 38.3 58.5 29.5 24.7 14.2 29.1 29.8
TroL-3.8B 22.5 65.3 70.2 63.0 69.7 20.3 39.8 50.1
TroL-7B 23.6 44.7 74.4 38.6 40.9 16.0 32.3 38.6
(e) Evaluating sub-benchmark in VisualWebBench Liu et al. (2024b) with numerous open-source LLVMs.
Table 3: Detailed comparison of Refer to captionTroL across challenging evaluation benchmarks. Note that, the sub-benchmark category names in (c) and (d) are represented in Appendix A. Note that, Phi-3-mini Abdin et al. (2024) built in TroL-3.8B has shown excellent benefits for understanding web pages and solving human-level math problems.
Implementation Detail.
To ensure successful reproducibility, we present four key technical aspects of Refer to captionTroL: the detailed structure of (a) backbone multimodal LLMs, (b) vision encoder, vision projectors, TroL Gating. In addition, the detailed procedures of (c) training and inference are described.
For backbone multimodal LLMs, we employ Phi-3-mini (Abdin et al., 2024) and InternLM2 (Team, 2023; Cai et al., 2024), where Phi-3-mini 3.8B model consists of 32 layers with hidden dimension of 3072, while InternLM2-1.8B | 7B features 24 | 32 layers with hidden dimension of 2048 | 4096, respectively.
We use CLIP-L (Radford et al., 2021) and InternViT (Chen et al., 2023c) as vision encoders, each comprising 428M | 300M parameters, with 24 layers and hidden dimension of 1024. When investigating best structural combination, we consider CLIP-L for 1.8B and 7B model and consider InternViT for 3.8B model. The vision projector consists of an MLP that adjusts the hidden dimension from 1024 to 2048 | 3072 | 4096 to match the hidden dimension of backbone multimodal LLMs. This MLP contains two fully-connected layers with GELU activation (Hendrycks and Gimpel, 2016). TroL Gating employs a single fully-connected layer that converts the hidden dimension from 2048 | 3072 | 4096 to 1, resulting in a total of $2048\times 24=$49K, $3072\times 32=$98K, and $4096\times 32=$131K parameters for TroL Gating each, which are minimal compared to the 1.8B, 3.8B, and 7B model sizes.
The training and evaluation of Refer to captionTroL are conducted in a computing environment with 8$\times$×NVIDIA Tesla A100 80GB and 8$\times$×NVIDIA RTX A6000 48GB, each. To optimize the training process, we use one epoch of training for each step under 4/8-bit quantization and bfloat16 data type (Kalamkar et al., 2019) for each backbone multimodal LLM: Refer to captionTroL-1.8B (4-bit), Refer to captionTroL-3.8B (8-bit), and Refer to captionTroL-7B (4-bit). The 4-bit quantization employs double quantization and normalized float 4-bit (nf4)(Dettmers et al., 2023). Additionally, QLoRA (Hu et al., 2021; Dettmers et al., 2023) is used to train the multimodal LLMs with 64 rank and 64 alpha parameters. We apply the AdamW optimizer (Loshchilov and Hutter, 2019) and use cosine annealing to schedule the learning rate from 1e-4 to 1e-6 in each training step. We also utilize gradient checkpointing (Sohoni et al., 2019) for efficient memory usage. With a gradient accumulation of 6, batch sizes are totally set to 768 for each training step, and each step takes approximately one to three days according to the model sizes. For efficient inference, we validate Refer to captionTroL under the same quantization bit during training, and we employ deterministic beam search (Freitag and Al-Onaizan, 2017) ($n=5$) for text generation. Moreover, in inference, we apply layer traversing technique only to the user questions in order to avoid dramatically increased inference time. Note that, Refer to captionTroL is implemented on the efficient propagation, FlashAttention2 (Dao et al., 2022; Dao, 2023) for speed-up attention computation.
LLMs Param MMStar MM-Vet
Gemma 2B 44.6 42.8
InternLM2 1.8B 45.5 45.1
Qwen1.5 4B 45.5 49.9
Phi-3-mini 3.8B 46.5 51.1
Mistral 7B 49.7 53.2
InternLM2 7B 51.3 54.7
(a) Backbone pre-trained LLMs
Operation MMStar MM-Vet
$L(x)\odot L(L(x))$ 23.5 28.8
$(1-w)\odot L(x)$ 40.8 45.4
$w\odot L(L(x))$ 44.3 48.1
Random $w$ 40.3 37.0
Uniform $w$ 41.4 46.3
Figure 5 51.3 54.7
(d) Mixing operation
Family Lay-Trav MMStar MM-Vet
TroL-1.8B 36.0 34.6
TroL-1.8B 45.5 45.1
TroL-3.8B 37.4 43.5
TroL-3.8B 46.5 51.1
TroL-7B 41.2 45.8
TroL-7B 51.3 54.7
(b) Use of layer traversing
Percent MMStar MM-Vet
10% 39.9 41.5
30% 46.3 48.2
50% 50.8 54.0
70% 51.2 54.4
90% 51.3 54.6
100% 51.3 54.7
(e) Training percentage
Structure Param MM-Vet MM-Vet
PR $\times$2 13B 52.3 57.4
PR 6B 52.2 57.1
MHA×$\times$2 + FC 5B 51.8 56.8
MHA + FC 2B 51.7 56.6
FC×$\times$2 537M 51.3 54.8
FC 131K 51.3 54.7
(c) TroL Gating
LLVMs VRAM Time Ratio
InternLM2-1.8B 6GB 1.0
+CLIP-L (Image Tokens) +1GB 1.1
+Layer Traversing +1GB 1.3
Phi-3-mini-3.8B 7GB 1.0
+InternViT (Image Tokens) +1GB 1.2
+Layer Traversing +1GB 1.4
(f) Inference speed
Table 4: Ablation studies to identify the effectiveness of Refer to captionTroL by controlling the six main factors.
Figure 6: Visualization of layer traversing activated degree in each layer where mixing ratios of $w$ in token-wise manner are shown as the colors in color bar. To clearly discriminate where the layer traversing mostly occurs, they are applied to min-max normalization in each layer. (i.e., The brighter they are, the more layer traversing happens)
Validation on Evaluation Benchmarks.
We have demonstrated super vision language performances of Refer to captionTroL, despite its limited model size, as depicted in Table 1, Table 2, and Table 4. This is attributed to the enhanced learning capabilities by layer traversing. Furthermore, Table 4 have shown several ablation studies to clearly corroborate the effectiveness of Refer to captionTroL in light of seven factors: (a) backbone pre-trained LLMs, (b) the use of layer traversing, (c) the structure of TroL Gating, (d) mixing operation of TroL-Mixer, (e) training percentage of visual instruction tuning, and (f) ratio of measured time for inference speed. Note that, in Table 4(c), ‘FC’ denotes a fully-connected layer, and ‘MHA’ denotes a multi-head attention block, and ‘PR’ denotes the structure of Perceiver Resampler (Alayrac et al., 2022). Through the ablation, we consider the efficiency of the model structure and performance on vision language tasks to build the current architectures of Refer to captionTroL, including the pre-trained LLM, TroL Gating architecture, and the TroL-Mixer operation. Additionally, layer traversing focusing on user’s questions during inference makes text generation speed comparable to the baselines we used for pre-trained LLM. Note that, all descriptions of the evaluation benchmarks used in this paper are explained in Appendix A, and we show diverse samples for Refer to captionTroL’s text generation quality in Appendix B.
5 Discussion and Conclusion
A new LLVM family, Refer to captionTroL, has demonstrated significant advancements in vision language performance despite its inherent limitation of having smaller layers compared to larger model sizes in both open- and closed-source LLVMs. Table 4 suggests that reusing layers through layer traversing can be an effective alternative to incorporating additional modules, such as mixed vision encoders and multiple computer vision models. We expect Refer to captionTroL to remain a competitive and efficient option in the rapidly evolving field of vision language models, solidifying its place in the landscape of efficient LLVMs.
Interestingly, when analyzing all mixing ratios for each layer, we observed in Figure 6 that the layer traversing event of looking back and retracing the answering stream mostly occurs in shallower layers, while the deeper layers are not involved in traversing. This suggests that recursively enhancing vision language features continues until they are fully matured; once matured, no need for anymore. Further, we plan to explore further methods in the shallower and deeper layers, which significantly enhance learning capabilities by virtually increasing the hidden dimension without physically enlarging it. We believe this approach, combined with layer traversing, could be one of the crucial keys serving as efficient LLVMs, potentially propelling Refer to captionTroL to the forefront within 1$\sim$3B.
6 Limitations
TroL, like several other LLVMs, naturally faces the challenge of the training computational burden. For building Refer to captionTroL, we used high-end GPUs (8$\times$×NVIDIA Tesla A100 80GB) and take at most three days in training despite using them. Practically, this challenge might be mitigated using numerous optimization tools and techniques (Xue et al., 2024), such as PagedAttention (Kwon et al., 2023), ChunkAttention (Ye et al., 2024), optimized CUDA kernel & HIP graph, GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023a), SqueezeLLM (Kim et al., 2023c), dynamic KV cache, FP8 KV cache, and prefix caching. However, these techniques mentioned tend to be more applicable during the inference phase. Therefore, advanced techniques beyond quantization for reducing the training computational burden should be further developed to enable the AI community to handle large models more effectively. Beyond that, we expect Refer to captionTroL to be equipped with numerous bootstrapping methods (Lee, 2020; Lee et al., 2021; Kim et al., 2021; Lee et al., 2022; Kim et al., 2023b; Lee et al., 2023; Kim et al., 2023a, d; Lee et al., 2024a; Park et al., 2024c, b; Kim et al., 2024), providing a wide range of variations for both general and specific tasks.
7 Ethics Statement
We affirm that all research presented in this paper adheres to the highest principles of ethical conduct and integrity. The experiments conducted and the results reported are based on rigorous scientific methodologies, with the goal of contributing positively and responsibly to the field of large language and vision models (LLVMs). All datasets utilized in this study, including visual instruction tuning datasets (Chen et al., 2023b, 2024b; Li et al., 2024b; Hu et al., 2024a; Gao et al., 2023; Wang et al., 2024a; Yue et al., 2023, 2024), were acquired and analyzed in strict compliance with relevant regulations and guidelines governing research ethics and data privacy. We ensured that all data handling processes protected the confidentiality and rights of individuals represented in the datasets. Furthermore, we have transparently discussed any potential limitations and ethical considerations in Section 6, Limitations. This commitment to transparency allows for a critical assessment of our work and underscores our dedication to maintaining the highest standards of integrity and accountability. We are committed to respecting and considering the impact of our research on communities and individuals. Our goal is to advance the field responsibly, with a conscientious approach that values ethical considerations as much as scientific innovation. By upholding these principles, we aim to foster trust and integrity within AI community.
Appendix A Evaluation Benchmarks
  • Q-Bench (Wu et al., 2023) is designed to evaluate the low-level visual abilities of Multi-modality Large Language Models (MLLMs). It is segmented into three primary categories: perception, description, and assessment. The perception section focuses on the ability of MLLMs to identify and interpret basic image attributes. The description section checks the precision and completeness of how MLLMs can articulate these attributes. The assessment section measures the extent to which MLLMs’ evaluations of image quality match human judgments. The dataset contains a total of 81,284 samples.
  • SQA-${}^{\text{I}}$) (Lu et al., 2022) is part of the broader ScienceQA (SQA) dataset, which aims to improve reasoning and interpretability in AI systems through science-based question answering. This dataset covers a wide range of science disciplines, featuring 26 different topics in natural, social, and language sciences, all accompanied by annotated answers, lectures, and explanations. SQA-IMG includes image-related samples, amounting to 10,332 question-answer pairs.
  • AI2D (Kembhavi et al., 2016) or AI2 Diagrams, addresses diagram interpretation and reasoning challenges, focusing on syntactic parsing and semantic understanding. It supports research into diagram structure and element relationships, critical for tasks like diagram-based question answering. This collection includes over 5,000 diagrams from elementary science topics, along with over 15,000 multiple-choice questions.
  • ChartQA (Masry et al., 2022) develops to challenge and improve question answering systems that deal with data visualizations like bar charts, line charts, and pie charts. This dataset tests systems on questions requiring arithmetic and logical reasoning and includes both human-generated and machine-created question-answer pairs. It comprises 32,719 samples in total.
  • SEED-${}^{\text{I}}$) (Li et al., 2023c), a subset of SEED-Bench, evaluates the generative comprehension skills of multimodal large language models (MLLMs) with a focus on spatial and temporal understanding. It offers several subsets mapped to 12 evaluation dimensions across image and video modalities, with SEED-IMG specifically concentrating on images.
  • POPE (Li et al., 2023e) introduces a method to systematically assess the tendency of LLVMs to falsely generate nonexistent objects in images. This method turns the evaluation into a binary classification task using polling questions, providing a fair and adaptable approach.
  • HallusionBench (HallB) (Liu et al., 2023a) is crafted to evaluate and explore visual illusions and knowledge hallucinations in large language and vision models (LLVMs). This benchmark uses carefully crafted example pairs to identify model failures, featuring diverse visual-question pairs including subsets focused on illusions, math, charts, tables, maps, and OCR. It includes 346 images and 1,129 questions.
  • MME (Fu et al., 2023) serves as a comprehensive evaluation framework for Multimodal Large Language Models (MLLMs), focusing on various perception and cognition tasks through 14 sub-tasks like coarse and fine-grained recognition, OCR, and commonsense reasoning. This benchmark aims to address existing evaluation gaps and ensures a thorough testing environment for MLLMs.
  • MathVista (Lu et al., 2023b) is an extensive benchmark designed to test visual-based mathematical reasoning in AI models. It integrates visual understanding in evaluating models’ abilities to solve math problems that involve visuals. The dataset consists of three subsets: IQTest, FunctionQA, and PaperQA, totaling 6,141 examples.
[…]
  • MM-Vet (Yu et al., 2023) is designed to systematically evaluate LMMs on complex tasks requiring multiple vision language (VL) capabilities. It tests recognition, knowledge, OCR, spatial awareness, language generation, and math, integrating these abilities into 16 different task combinations. The dataset includes 200 images and 218 questions, each requiring the integration of multiple capabilities.
  • LLaVA Bench in the ${}^{\text{W}}$) (Liu et al., 2023c) assesses large multimodal models (LMM) on complex tasks and new domains through a collection of 24 images with 60 questions. This dataset features diverse settings, including indoor, outdoor, artworks, and memes, with each image accompanied by detailed descriptions and curated questions.
  • MMStar (Chen et al., 2024c) is crafted to precisely evaluate the true multimodal capabilities of LLVMs by ensuring that each sample critically relies on visual content for accurate answers while minimizing data leakage. It comprises 1,500 meticulously selected samples and is organized into six primary sub-benchmarks as follows:
    • Coarse perception (CP), which pertains to the ability to grasp and interpret the overarching features and themes of an image without focusing on minute details,
    • Fine-grained perception (FP), which denotes a detailed level of image comprehension that emphasizes the intricate and nuanced aspects of visual content,
    • Instance reasoning (IR), which encompasses advanced cognitive abilities aimed at understanding and interpreting individual and collective object attributes and their interrelations within an image,
    • Logical reasoning (LR), which involves a sophisticated framework of cognitive processes designed to interpret, deduce, and infer conclusions from visual content through a structured approach to logic and reasoning,
    • Science & technology (ST), which includes a comprehensive framework for the application and integration of knowledge across a wide range of scientific and technological domains,
    • Math (MA), which is a fundamental pillar of logical and analytical reasoning and includes a broad spectrum of skills essential for understanding, applying, and interpreting quantitative and spatial information.
  • MathVerse (Zhang et al., 2024) assesses the capabilities of Multi-modal Large Language Models (MLLMs) in visual mathematical reasoning, particularly their ability to understand visual diagrams and mathematical expressions. This dataset is categorized into three primary areas: plane geometry, solid geometry, and functions, and further detailed into twelve types like length and area, encompassing 2,612 visual mathematical challenges.
    To investigate how MLLMs process visual diagrams in mathematical reasoning, the creators of MathVerse developed six distinct versions of each problem, each version presenting different levels of multi-modal information. They initially established three specific classifications for the text content within the problems:
    • Descriptive Information, which includes content that is directly visible and explicitly depicted in the diagrams,
    • Implicit Property, which encompasses details that demand a more advanced visual perception yet less mathematical knowledge to interpret from the diagram,
    • Essential Condition, which pertains to crucial numerical or algebraic data necessary for solving the problem that cannot be inferred solely from the visual diagram.
    Based on these categories, to thoroughly assess the true visual understanding capabilities of MLLMs and their utility in multi-modal mathematical contexts, the researchers created six versions or sub-benchmarks of each problem in MathVerse, described as follows:
    • Text dominant (TD) version, which preserves all textual elements, including the three textual categories and the main question, prompting MLLMs to primarily depend on textual information.
    • Text lite (TL) version reduces the Descriptive Information from the Text dominant version, promoting reliance on the diagram for elementary data.
    • Text only (TO) version removes the visual elements entirely, focusing on textual content to discern where MLLMs predominantly derive contextual information for problem solving.
    • Vision intensive (VI) further excludes Implicit Property from the Text lite version, urging MLLMs to intensify their visual analysis to gather essential cues for mathematical reasoning.
    • Vision dominant (VD), evolving from the Text lite version, omits Essential Condition from the textual information and instead visually annotates these details in diagrams, compelling MLLMs to identify and accurately link these essential conditions solely through visual examination.
    • Vision only (VO) eliminates all textual descriptions, presenting the problem exclusively through visual means and challenging MLLMs to decode and identify mathematical queries purely based on visual data, serving as the ultimate test of their visual reasoning skills in mathematics.
  • VisualWebBench (Liu et al., 2024b) assesses the capabilities of multimodal large language models (MLLMs) specifically in the web domain. It is designed to address the lack of a comprehensive benchmark that evaluates the unique characteristics of web pages and measures fine-grained abilities such as OCR, understanding, and grounding in text-rich and interactive web environments. It covers a wide range of domains, including science, travel, sports, engineering, and government, and tasks such as captioning, WebQA, heading OCR, element grounding, and action prediction, containing a total of 1,534 instances.
Appendix B Refer to captionTroL Generation Quality